# Authors’ Instructions

### v3

@@ Abstract
As the cloud infrastructure grows, it becomes more challenging to manage resources in such a massive, diverse, and distributed setting, despite the fact that cloud computing provides computational capabilities on-demand. Due to resource variability and unpredictability, resource allocation issues arise in a cloud setting. A Quality of Service (QoS) based autonomic resource management strategy automates resource management, delivering trustworthy, dependable, and cost-effective cloud services that efficiently execute workloads. Autonomic cloud computing aims to understand how computing systems may autonomously accomplish user-specified ”control” objectives without the need for an administrator and without violating the Service Level Agreement (SLA) in a dynamic cloud computing environments. This article presents a research perspective and analysis on autonomous resource allocation in cloud computing, with a focus on QoS and SLA-aware autonomous resource management. The study also discusses the current status of autonomic resource management in the cloud and highlights key next-generation research directions.
@@
Cloud Computing, Autonomic Cloud Computing, Autonomic Computing, Service Level Agreement, Quality of Service
@@ 1 Introduction
Cloud computing is an evolving utility computing mechanism in which cloud consumers can detect, choose, and utilize the resources (infrastructure, software and platform) and provide service to users based on a pay-per-use model as computing utilities [1]. The current computing mechanism is effective, particularly for medium and small cloud based companies, in that it permits easy and reliable access to cloud services like infrastructure, software, and platforms. Current cloud computing bears a strong resemblance to the established models of cluster computing and grid computing [2]. The important key technical features of cloud computing include autonomic service, rapid elasticity, end-to-end virtualization support, on-demand resource pooling, and transparency in cloud billing. Further, non-technical features of cloud computing include environment friendliness, little overhead maintenance, lower upfront costs, faster deployment times, Service Level Agreement (SLA) and pay-as-you-go- models [3]. In a distributed computing environment, the unpredictability of services is a fact, so the same possible in the cloud as well. The success of next-generation cloud computing infrastructures will depend on how capably these infrastructures will discover and dynamically tolerate computing platforms, that meet the randomly varying resource and service requirements of cloud customer applications [4].
@@ 2 Service Level Agreements (SLAs)
SLA designates what you require from your consumers/service customers in order to provide the service specified. It needs assurance and support from both parties to provision and follow the contract in order for the SLA to work efficiently [5]. In the SLA, both parties (the cloud provider and the cloud consumer) should have specified the possible deviations to achieve appropriate quality attributes. In other words, it is vital for users to acquire assurances from suppliers on service provisions. Typically, providers and customers negotiate SLAs to deliver these assurances. The very first problem is the description of SLA terms in such a way that has a suitable level of granularity, namely the compromises between accuracy and complexity, so that they can ensure most of the user hopes and are comparatively simple to be prejudiced, certified, calculated, and imposed by the resource provisioning mechanism on the cloud [3]. In addition, different cloud service models (IaaS (infrastructure as a service), PaaS (platform as a service), and SaaS (software as a service)) will need to express different SLA meta disclaimers [6]. This also increases the number of implementation issues for cloud providers. Moreover, innovative SLA mechanisms require continuous integration of consumer response and customization features into the SLA assessment framework. As cloud service models develop and become omnipresent, it increases the probability of clarifying the way the services are provisioned and managed. As a result, it allows providers to address the various needs of their customers [7]. In this perspective, SLAs appear as a significant characteristic that subsequently serves as the basis for the predictable quality level of the services made available to customers by the providers. Nonetheless, the collection of the recommended SLAs by providers (with marginal overlaps), has led to manifold different definitions of cloud SLAs [8]. Moreover, confusion exists as to what the difference is between SLAs and agreements, what the marginal quality is, what terms are involved in each one of these documents, and if and how they are associated.
@@ 3 SLA and QoS in Cloud
Logically, based on QoS requirements such as scalability, high availability, trust, and security, these applications will be characterized, identified in the so-called SLAs [4]. Figure 1 shows the process of SLA negotiation between cloud users and cloud providers. The current cloud technology is not completely personalized to honor probable SLAs, though both industrial and academic research groups are presenting increasing interest in problems of QoS assurance within the context of cloud computing [9]. In general, an SLA necessitates a precise assessment of the characteristics of the required resources. Application services introduced in clouds (e.g., web applications, web services) are frequently characterized by great load inconsistency; therefore, the amount of resources required to honor their SLAs may vary, particularly over time. An important challenge for cloud providers is to automate the management of virtual servers while keeping into account both the high-level QoS requirements of hosted applications and resource supervision expenses [10]. Cloud market mechanisms are consistently static and cannot react to dynamic variations in consumer desires. To respond to these issues, there is a requirement for an adaptive methodology for autonomically springing SLA patterns based on consumer requirements. The present research in cloud SLA limits the capability of matching conformation metrics to acceptable. These metrics comprise statistical measures such as the standard deviation that want to be computed from the expected and actual outcomes of services delivered to customers [11]. Semantic Web technologies can be used to improve the descriptions and, therefore, increase the quality of these matches. Although cloud consumers do not have full supervisory control over the fundamental computing resources, they do require ensuring attributes such as quality, accessibility, trustworthiness, and performance of these resources when users have transferred their fundamental business functions onto their honoured cloud [12].
Figure 1: SLA Negotiation between Cloud Users and Cloud Providers
Figure 2: Comparison of SLA deviations
@@ 4 Analysis of SLA Deviation
Table 1: Cloud SLA Deviations
@@
Figure 3: Actual SLA Deviation
@@ 5 Existing Challenges in Traditional Cloud Computing
Usually, cloud providers do not consider compensation because resource providers do not actually enforce penalties for SLA violations. There should be a penalty delay cost or consumers’ compensation if the cloud provider misses the deadline [13]. Furthermore, when the cloud provider violates the terms, it transfers the risk to IaaS providers. Penalty delay cost is how much the service provider has to give concessions to users for SLA violations. It is dependent on the penalty rate and the penalty delay time period [14]. Researchers have tried to reduce the effect of inaccuracy using two approaches: first, considering the penalty compensation clause in SLAs with IaaS providers and imposing SLA violations; and second, adding some slack time during scheduling to avoid risk [15]. However, this increases the cost and execution time significantly [11] [12]. We can use autonomous cloud computing to solve this problem, understanding how computing systems can autonomously accomplish user-specified ”control” objectives without the need for an administrator, without compromising QoS, and without violating the SLA in dynamic cloud computing environments.
@@ 6 Potential Solution: Autonomic Cloud Computing
The Autonomic Computing Initiatives (ACI) at IBM became one of the earliest sector-wide programmes to focus on designing computers to work well with no or minimal intervention by humans [16]. IBM’s Tivoli Frameworks group originally tested autonomic computing to optimise DB2 efficiency. The endeavour had an enormous impact on studies of how the nervous systems of humans work together. For example, how the autonomic nervous system responds to stimulation even when the user isn’t consciously aware of it and how an autonomic computing environment uses advanced computational intelligence without the user even realising it’s happening [17]. Additionally, humans’ nervous systems can do many simple things at the same time, like change the internal environment, change the rate of breathing, and release chemicals through pores in response to stimuli. All of this happens while following rules and ”limits” and based on signals we receive or sense from our surroundings or from within ourselves [16]. An autonomous technology ecosystem operates on data collected, sensed, or taught, akin to how the human body processes information, without requiring immediate human oversight over system management operations.
@@ 6.1 Background
Biological networks are a source of inspiration for QoS-based autonomous systems because of their remarkable ability to deal with unpredictability, variability, change, errors, and so forth. Executing a programme within a deadline while satisfying user-described QoS standards with minimal complexity is the purpose of autonomic systems [17]. The human’s autonomic nervous system (ANS) is a primary model for autonomous computer systems. When faced with a volatile environment, the ANS can dynamically handle and manage any circumstance that may arise [23]. Similar to how the ANS regulates bodily activities like respiration and digestion, autonomous cloud computing systems (ACCSs) automate the operation of cloud-based software and systems. In the same way that the ANS checks and monitors, ACCSs react situationally by doing things like self-optimizing, self-configuring, self-protecting, and self-healing [18].
6.2 Properties
Autonomic computing, also called self-adaptation mechanisms, is an area of study that focuses on how computers may independently accomplish preferred actions. ”Self-*” systems are a popular way to describe these kinds of systems; the ”*” indicates the sort of actions, such as Self-”configuration, optimization, protection, healing” [19]. Figure 4 shows the four properties of autonomic computing system.
Figure 4: Properties of Autonomic Computing System
A self-configuring autonomic system is one that can adjust to new conditions without external interference [22]. Based on error signals or warnings issued by a monitoring device, the framework repairs faulty or outdated elements autonomously [20]. Autonomic systems that can optimise themselves by decreasing resource overload and under-utilisation and effectively performing computing tasks are called self-optimizing [18]. An autonomic system’s ability to self-protect means it is capable of warding off intruders and hackers [19]. The system’s autonomic controller must additionally be able to recognise and avoid damaging attacks. An autonomous system can self-diagnose, self-evaluate, and self-repair when it makes mistakes; no human interaction is required for this process, referred to as self-healing [22]. This self-* attribute improves performance by reducing the impact of errors on execution, a feature known as fault tolerance. In an ideal world, self-managing and self-healing systems would never require human intervention for setup or maintenance. When taken as a whole, the above-mentioned characteristics must fall under the supervision of self-managed systems [21].
@@ 6.3 Architecture
Various real-world systems accomplish these goals with varied degrees of precision and effectiveness. The amount of oversight and involvement from humans can also differ. The Autonomic Manager (AM) is an intelligent unit that operates under IBM’s Autonomic Computing architecture. It communicates with its surroundings through administrative graphical user interfaces such as sensors and effectors and takes action according to data gathered from sensors and policies stored in a low-level database. Using broad-brush notifications and directives, an administrator sets up the AM. IBM implements the autonomic process, as illustrated in Figure 5. Autonomic systems are frequently organised as monitor, analyse, plan, and execute (MAPE) phases.
Figure 5: Architecture of Autonomic Computing System
First, monitors interact with external hardware to obtain sensor data, which they then send to the next component for further analysis in order to regularly review QoS measures [23]. The monitoring module’s data is evaluated in the Analyse and Plan modules, which then formulate suitable responses to system alarms. This autonomic system responds appropriately to the alerts issued based on the data analysis results. The execution plan has to be put into effect by the executor, and its key purpose is to preserve the QoS of an execution programme after a comprehensive assessment involving verifying and validating it to offer assurance that the modification will truly operate [24]. An executor monitors the addition of new information to the database and responds to the results.
@@ 6.4 Validation
Based on QoS characteristics, smart autonomic systems take cues from biological systems that have proven proficient at dealing with challenges including dynamism, diversity, and the unpredictability of cloud resources. A smart autonomous framework aims to deploy applications on time while meeting the user-described QoS requirements with minimal complexity. The MAPE-K loop is the basis of an autonomous computing system, which takes into account all four phases of an autonomic system with a control loop: monitoring, analysing, planning, and execution. The simulation-based experiments have been conducted using the CloudSim toolkit [25] to verify the use of autonomic computing for cloud resource management, as demonstrated in notable previous research [19] [20] [21] [22] [23] [24]. Autonomic systems outperformed non-autonomic computing systems experimentally in terms of QoS (execution time) and SLA violation rate, as shown in Figure 6 and Figure 7, respectively, because the service can self-manage according to its environment’s demands. Through the implementation of autonomic computing, application utilisation and efficiency were maximised, while performance and other quality of service assurances were ensured [20]. These results prove without a reasonable doubt that autonomous cloud computing maintains system stability in the face of uncertainty and allows for rapid adaptation to novel environmental situations, such as software or hardware faults [21].
Figure 6: Comparison of Autonomic and Non-Autonomic Computing Systems based on QoS (execution time)
Figure 7: Comparison of Autonomic and Non-Autonomic Computing Systems based on SLA violation rate
6.5 AI for Autonomic Computing: Promising Future Directions
@@
Using data obtained on system functions, Artificial Intelligence (AI) and Machine Learning (ML) could potentially be employed to assist with and generate autonomic behaviours. For instance, ML methods may be employed to identify trends within the workload, which can then be utilised to enhance the administration of available resources [26]. The autonomic manager might also automatically call upon ML-based, continually changing system recognition approaches such as neural networks to do self-learning, which would help reduce the modelling uncertainty. This allows for the generation of both black-box and grey-box models of the controlled systems throughout a concept movement, which can then be checked for sanity or, worse, for the detection of extremely important changes to the system’s functioning. Autonomic systems are frequently organised as MAPE phases, and AI may be used in conjunction with management approaches to aid in the analysis and planning parts of these phases [28]. Important advantages for autonomic self-management come from combining feedback control with data-driven model creation using ML, which provides important advantages for autonomic self-management.
@@
Several computing frameworks, including cloud, fog, edge, serverless computing, and quantum technology, have included autonomous computing using AI and ML methods [29]. Whenever a system has many possible settings, autonomic computing approaches become very important. Optimising search across a larger number of possible alternatives is achievable when there is a larger potential range of parameters in which preferences can change. The best use case for autonomous computing approaches is in the background, as a programmable gateway that applications may directly call [30].
@@
Due to the inherent autonomy of most Peer-to-Peer (P2P) networks, there are several apps that can handle node failures, network setup and updates, and, to a lesser extent, optimise performance independently. More and more online services and data centre management tools are incorporating AI and ML-based self-management abilities [31]. This allows these platforms to swiftly adapt to changing demands. Schedulers and workflow administrators don’t always have autonomic characteristics because they can’t keep tabs on the system’s health and offer real-time feedback, which makes them less autonomous [32]. One way to increase these systems’ capabilities is to incorporate ”tuning” capabilities that utilise AI/ML approaches. Hadoop and MapReduce are two examples of self-managed computing systems that offer self-organization and self-healing features that make extensive resource utilisation possible.

### v4

# Autonomic Cloud Computing: Research Perspective

@@ Abstract
As the cloud infrastructure grows, it becomes more challenging to manage resources in such a massive, diverse, and distributed setting, despite the fact that cloud computing provides computational capabilities on-demand. Due to resource variability and unpredictability, resource allocation issues arise in a cloud setting. A Quality of Service (QoS) based autonomic resource management strategy automates resource management, delivering trustworthy, dependable, and cost-effective cloud services that efficiently execute workloads. Autonomic cloud computing aims to understand how computing systems may autonomously accomplish user-specified ”control” objectives without the need for an administrator and without violating the Service Level Agreement (SLA) in a dynamic cloud computing environments. This chapter presents a research perspective and analysis on autonomic resource allocation in cloud computing based on the last decade of conducted research with a focus on QoS and SLA-aware autonomic resource management. This study delves into the current state of autonomic resource management in the cloud and introduces a conceptual model for Artificial Intelligence (AI)-driven autonomic cloud computing. This model aims to optimise server load distribution and energy consumption, thus enhancing cost savings and environmental impact. Finally, it highlights key next-generation research directions.
@@
Cloud Computing, Autonomic Cloud Computing, Artificial Intelligence, Autonomic Computing, Service Level Agreement, Quality of Service
@@ 1 Introduction
Cloud computing is an evolving utility computing mechanism in which cloud consumers can detect, choose, and utilize the resources (infrastructure, software and platform) and provide service to users based on a pay-per-use model as computing utilities A1 . The current computing mechanism is effective, particularly for medium and small cloud based companies, in that it permits easy and reliable access to cloud services like infrastructure, software, and platforms. Current cloud computing bears a strong resemblance to the established models of cluster computing and grid computing A2 . The important key technical features of cloud computing include autonomic service, rapid elasticity, end-to-end virtualization support, on-demand resource pooling, and transparency in cloud billing. Further, non-technical features of cloud computing include environment friendliness, little overhead maintenance, lower upfront costs, faster deployment times, Service Level Agreement (SLA) and pay-as-you-go- models A3 . In a distributed computing environment, the unpredictability of services is a fact, so the same possible in the cloud as well. The success of next-generation cloud computing infrastructures will depend on how capably these infrastructures will discover and dynamically tolerate computing platforms, that meet the randomly varying resource and service requirements of cloud customer applications A4 .
@@ 2 Service Level Agreements (SLAs)
SLA designates what you require from your consumers/service customers in order to provide the service specified. It needs assurance and support from both parties to provision and follow the contract in order for the SLA to work efficiently A5 . In the SLA, both parties (the cloud provider and the cloud consumer) should have specified the possible deviations to achieve appropriate quality attributes. In other words, it is vital for users to acquire assurances from suppliers on service provisions. Typically, providers and customers negotiate SLAs to deliver these assurances. The very first problem is the description of SLA terms in such a way that has a suitable level of granularity, namely the compromises between accuracy and complexity, so that they can ensure most of the user hopes and are comparatively simple to be prejudiced, certified, calculated, and imposed by the resource provisioning mechanism on the cloud A3 . In addition, different cloud service models (IaaS (infrastructure as a service), PaaS (platform as a service), and SaaS (software as a service)) will need to express different SLA meta disclaimers A6 . This also increases the number of implementation issues for cloud providers. Moreover, innovative SLA mechanisms require continuous integration of consumer response and customization features into the SLA assessment framework. As cloud service models develop and become omnipresent, it increases the probability of clarifying the way the services are provisioned and managed. As a result, it allows providers to address the various needs of their customers A7 . In this perspective, SLAs appear as a significant characteristic that subsequently serves as the basis for the predictable quality level of the services made available to customers by the providers. Nonetheless, the collection of the recommended SLAs by providers (with marginal overlaps), has led to manifold different definitions of cloud SLAs A8 . Moreover, confusion exists as to what the difference is between SLAs and agreements, what the marginal quality is, what terms are involved in each one of these documents, and if and how they are associated.
Figure 1: SLA Negotiation between Cloud Users and Cloud Providers A23 A38
@@ 3 SLA and QoS in Cloud
Logically, based on QoS requirements such as scalability, high availability, trust, and security, these applications will be characterized, identified in the so-called SLAs A4 . Figure 1 shows the process of SLA negotiation between cloud users and cloud providers. The current cloud technology is not completely personalized to honor probable SLAs, though both industrial and academic research groups are presenting increasing interest in problems of QoS assurance within the context of cloud computing A9 . In general, an SLA necessitates a precise assessment of the characteristics of the required resources. Application services introduced in clouds (e.g., web applications, web services) are frequently characterized by great load inconsistency; therefore, the amount of resources required to honor their SLAs may vary, particularly over time. An important challenge for cloud providers is to automate the management of virtual servers while keeping into account both the high-level QoS requirements of hosted applications and resource supervision expenses A10 . Cloud market mechanisms are consistently static and cannot react to dynamic variations in consumer desires. To respond to these issues, there is a requirement for an adaptive methodology for autonomically springing SLA patterns based on consumer requirements. The present research in cloud SLA limits the capability of matching conformation metrics to acceptable. These metrics comprise statistical measures such as the standard deviation that want to be computed from the expected and actual outcomes of services delivered to customers A11 . Semantic Web technologies can be used to improve the descriptions and, therefore, increase the quality of these matches. Although cloud consumers do not have full supervisory control over the fundamental computing resources, they do require ensuring attributes such as quality, accessibility, trustworthiness, and performance of these resources when users have transferred their fundamental business functions onto their honoured cloud A12 .
Figure 2: Comparison of SLA deviations A20 A23
@@ 4 Analysis of SLA Deviation
Table 1: Cloud SLA Deviations A20 A23
@@
Figure 3: Actual SLA Deviation A20 A23 A23
@@ 5 Existing Challenges in Traditional Cloud Computing
Usually, cloud providers do not consider compensation because resource providers do not actually enforce penalties for SLA violations. There should be a penalty delay cost or consumers’ compensation if the cloud provider misses the deadline A13 . Furthermore, when the cloud provider violates the terms, it transfers the risk to IaaS providers. Penalty delay cost is how much the service provider has to give concessions to users for SLA violations. It is dependent on the penalty rate and the penalty delay time period A14 . Researchers have tried to reduce the effect of inaccuracy using two approaches: first, considering the penalty compensation clause in SLAs with IaaS providers and imposing SLA violations; and second, adding some slack time during scheduling to avoid risk A15 . However, this increases the cost and execution time significantly A11 A12 . Autonomic cloud computing can be utilized to solve this problem, understanding how computing systems can autonomously accomplish user-specified ”control” objectives without the need for an administrator, without compromising QoS, and without violating the SLA in dynamic cloud computing environments.
@@ 6 Potential Solution: Autonomic Cloud Computing
The Autonomic Computing Initiatives (ACI) at IBM became one of the earliest sector-wide programmes to focus on designing computers to work well with no or minimal intervention by humans A16 . IBM’s Tivoli Frameworks group originally tested autonomic computing to optimise DB2 efficiency. The endeavour had an enormous impact on studies of how the nervous systems of humans work together. For example, how the autonomic nervous system responds to stimulation even when the user isn’t consciously aware of it and how an autonomic computing environment uses advanced computational intelligence without the user even realising it’s happening A17 . Additionally, humans’ nervous systems can do many simple things at the same time, like change the internal environment, change the rate of breathing, and release chemicals through pores in response to stimuli. All of this happens while following rules and ”limits” and based on signals we receive or sense from our surroundings or from within ourselves A16 . An autonomous technology ecosystem operates on data collected, sensed, or taught, akin to how the human body processes information, without requiring immediate human oversight over system management operations.
@@ 6.1 Background
Biological networks are a source of inspiration for QoS-based autonomous systems because of their remarkable ability to deal with unpredictability, variability, change, errors, and so on. Executing a programme within a deadline while satisfying user-described QoS standards with minimal complexity is the purpose of autonomic systems A17 . The human’s autonomic nervous system (ANS) is a primary model for autonomic computing systems. When faced with a volatile environment, the ANS can dynamically handle and manage any circumstance that may arise A23 . Similar to how the ANS regulates bodily activities like respiration and digestion, autonomic cloud computing systems (ACCSs) automate the operation of cloud-based software and systems. In the same way that the ANS checks and monitors, ACCSs react situationally by doing things such as self-optimizing, self-configuring, self-protecting, and self-healing A18 .
6.2 Self-* Properties
Autonomic computing, also called self-adaptation mechanisms, is an area of study that focuses on how computers may independently accomplish preferred actions. ”Self-*” systems are a popular way to describe these kinds of systems; the ”*” indicates the sort of actions, such as Self-”configuration, optimization, protection, healing” A19 . Figure 4 shows the four properties of autonomic computing system.
Figure 4: Self-* Properties of Autonomic Computing System A16 A38 A23
A self-configuring autonomic system is one that can adjust to new conditions without external interference A22 . Based on error signals or warnings issued by a monitoring device, the framework repairs faulty or outdated elements autonomously A20 . Autonomic systems that can optimise themselves by decreasing resource overload and under-utilisation and effectively performing computing tasks are called self-optimizing A18 . An autonomic system’s ability to self-protect means it is capable of warding off intruders and hackers A19 . The system’s autonomic controller must additionally be able to recognise and avoid damaging attacks. An autonomic system can self-diagnose, self-evaluate, and self-repair when it makes mistakes; no human interaction is required for this process, referred to as self-healing A22 . This self-* attribute improves performance by reducing the impact of errors on execution, a feature known as fault tolerance. In an real world, self-managing and self-healing systems would never require human intervention for setup or maintenance. When taken as a whole, the above-mentioned characteristics must fall under the supervision of self-managed systems A21 .
@@ 6.3 Architecture
Various real-world systems accomplish these goals with varied degrees of precision and effectiveness. The amount of oversight and involvement from humans can also differ. The Autonomic Manager (AM) is an intelligent unit that operates under IBM’s Autonomic Computing architecture A16 . It communicates with its surroundings through administrative graphical user interfaces such as sensors and effectors and takes action according to data gathered from sensors and policies stored in a low-level database. Using broad-brush notifications and directives, an administrator sets up the AM. IBM implements the autonomic process, as illustrated in Figure 5. Autonomic systems are frequently organised as monitor, analyse, plan, and execute (MAPE) phases.
Figure 5: General Architecture of Autonomic Computing System A16 A38 A23
First, monitors interact with external hardware to obtain sensor data, which they then send to the next component for further analysis in order to regularly review QoS measures A23 . The monitoring module’s data is evaluated in the Analyse and Plan modules, which then formulate suitable responses to system alarms. This autonomic system responds appropriately to the alerts issued based on the data analysis results. The execution plan has to be put into effect by the executor, and its key purpose is to preserve the QoS of an execution programme after a comprehensive assessment involving verifying and validating it to offer assurance that the modification will truly operate A24 . An executor monitors the addition of new information to the database and responds to the results.
Figure 6: Comparison of Autonomic and Non-Autonomic Computing Systems based on QoS (execution time) A20 A24 A40
Figure 7: Comparison of Autonomic and Non-Autonomic Computing Systems based on SLA violation rate A20 A24 A40
@@ 6.4 Validation
Based on QoS characteristics, smart autonomic systems take cues from biological systems that have proven proficient at dealing with challenges the include dynamism, diversity, and the unpredictability of cloud resources. A smart autonomous framework aims to deploy applications on time while meeting the user-described QoS requirements with minimal complexity. The MAPE-K loop is the basis of an autonomic computing system, which takes into account all four phases of an autonomic system with a control loop: monitoring, analysing, planning, and execution A16 . The simulation-based experiments have been conducted using the CloudSim toolkit A25 to verify the use of autonomic computing for cloud resource management, as demonstrated in notable previous research A19 A20 A21 A22 A23 A24 . Autonomic systems outperformed non-autonomic computing systems experimentally in terms of QoS (execution time) and SLA violation rate, as shown in Figure 6 and Figure 7, respectively, because the service can self-manage according to its environment’s demands. Through the implementation of autonomic computing, application utilisation and efficiency were maximised, while performance and other quality of service assurances were ensured A20 . These results prove without a reasonable doubt that autonomic cloud computing maintains system stability in the face of uncertainty and allows for rapid adaptation to novel environmental situations, such as software or hardware faults A21 .
7 AI and Computing: Intertwined in Autonomic Cloud Computing
@@
Using data obtained on system functions, Artificial Intelligence (AI) and Machine Learning (ML) could potentially be employed to assist with and generate autonomic behaviours. For instance, ML methods may be employed to identify trends within the workload, which can then be utilised to enhance the administration of available resources A26 . The autonomic manager might also automatically call upon ML-based, continually changing system recognition approaches such as neural networks to do self-learning, which would help reduce modelling uncertainty. This allows for the generation of both black-box and grey-box models of the controlled systems throughout a concept movement, which can then be checked for sanity or, worse, for the detection of extremely important changes to the system’s functioning. Autonomic systems are frequently organised as MAPE phases, and AI may be used in conjunction with management approaches to aid in the analysis and planning parts of these phases A27 . Important advantages of autonomic self-management come from combining feedback control with data-driven model creation using ML, which provides important advantages for autonomic self-management.
@@
Several computing frameworks, including cloud, fog, edge, serverless computing, and quantum technology, have included autonomic computing using AI and ML methods A28 . Whenever a system has many possible settings, autonomic computing approaches become very important. Optimising search across a larger number of possible alternatives is achievable when there is a larger potential range of parameters in which preferences can change. The best use case for autonomic computing approaches is in the background, as a programmable gateway that applications can directly call A29 .
@@
Due to the inherent autonomy of most Peer-to-Peer (P2P) networks, there are several apps that can handle node failures, network setup and updates, and, to a lesser extent, optimise performance independently. More and more online services and data centre management tools are incorporating AI and ML-based self-management abilities A30 . This allows these platforms to swiftly adapt to changing demands. Schedulers and workflow administrators don’t always have autonomic characteristics because they can’t keep tabs on the system’s health and offer real-time feedback, which makes them less autonomous A31 . One way to increase these systems’ capabilities is to incorporate ”tuning” capabilities that utilise AI/ML approaches. Hadoop and MapReduce are two examples of self-managed computing systems that offer self-organization and self-healing features that make extensive resource utilisation possible.
7.1 Transformative Effects of AI on Autonomic Cloud Computing